OpenStudy (anonymous):

You get 25% discount off a $465 TV as a special. The tax is 9%. How much is your final payment?

OpenStudy (ranga):

25% discount means you pay 75% of the cost. Find how much the TV costs by finding 75% of $465. Then, add tax. Tax = 9% means you multiply by 1.09 to find the total cost with tax included.

OpenStudy (anonymous):

i think i get it so i multiply the cost of the tv by 0.75 and multiply that by 1.09... right?

OpenStudy (anonymous):

This is how you solve this. remember this \[\frac{ is }{ off }\times \frac{ percent }{100 }\]

OpenStudy (ranga):

correct.

OpenStudy (anonymous):

thanks guys

OpenStudy (ranga):

you are welcome.